Quantitative Aptitude प्रश्न और उत्तर का अभ्यास करें

प्र: A train covers 180 km distance in 5 hours. Another train covers the same distance in 1 hour less. What is the difference in the distances covered by these trains in one hour if they are moving in the same direction? 2320 0

  • 1
    15 kms
    सही
    गलत
  • 2
    9 kms
    सही
    गलत
  • 3
    6 kms
    सही
    गलत
  • 4
    18 kms
    सही
    गलत
  • उत्तर देखेंउत्तर छिपाएं
  • Workspace

उत्तर : 2. "9 kms"
व्याख्या :

Answer: B) 9 kms Explanation: The first train covers 180 kms in 5 hrs => Speed = 180/5 = 36 kmph Now the second train covers the same distance in 1 hour less than the first train => 4 hrs => Speed of the second train = 180/4 = 45 kmph Now, required difference in distance in 1 hour = 45 - 36 = 9 kms.

प्र: 14 persons are seated around a circular table. Find the probability that 3 particular persons always seated together. 2318 0

  • 1
    11/379
    सही
    गलत
  • 2
    21/628
    सही
    गलत
  • 3
    24/625
    सही
    गलत
  • 4
    26/247
    सही
    गलत
  • उत्तर देखेंउत्तर छिपाएं
  • Workspace

उत्तर : 3. "24/625"
व्याख्या :

Answer: C) 24/625 Explanation: Total no of ways = (14 – 1)! = 13! Number of favorable ways = (12 – 1)! = 11!   So, required probability = 11!×3!13! = 39916800×66227020800 = 24625

प्र: Mr. X sold two properties P1 and P2 for Rs 1,00,000 each. He sold property P1 for 20% less than what he paid for it. What is the percentage of profit of property P2, so that he is not in gain or loss on the sale of two properties? 2316 0

  • 1
    33.33%
    सही
    गलत
  • 2
    29.97%
    सही
    गलत
  • 3
    25%
    सही
    गलत
  • 4
    22.22%
    सही
    गलत
  • उत्तर देखेंउत्तर छिपाएं
  • Workspace

उत्तर : 1. "33.33%"
व्याख्या :

Answer: A) 33.33% Explanation: Let 'A' be the cost price of property P1. Then from the given data, the selling price of P1 = Rs. 1,00,000 He got 20% loss on selling P1 ⇒A - 20A100 = 100000⇒A = 1,25,000 Therefore, the amount he lossed on selling P1 = 25,000 As ge he got no loss or gain on sale of P1 and P2, the gain on selling P2 = 25,000 But the selling price of P2 = 1,00,000 => Cost price of P2 = 75,000   Hence, the profit percentage on P2 = GainCPx 100 = 2500075000x100 = 1003= 33.33%

प्र: In a party there are male and female members. If 15 female quit then the number of females will become double the number of males. If 45 males quit no. of female becomes five times the number of males. Find the number of females ? 2315 0

  • 1
    175
    सही
    गलत
  • 2
    145
    सही
    गलत
  • 3
    165
    सही
    गलत
  • 4
    135
    सही
    गलत
  • उत्तर देखेंउत्तर छिपाएं
  • Workspace

उत्तर : 1. "175"
व्याख्या :

Answer: A) 175 Explanation: Let the number of males be 'M'Let the number of females be 'F'From the given conditions,(F-15)=2M.....(i)F=5(M-45).....(ii)By solving the equations (1) & (2) we get M = 80 Then substitute M = 80 in (1)F = 2M + 15F = 2 x 80 + 15=> F = 175Therefore, number of males = 80 and females = 175 in the party.

प्र: Find the odd one in the given series ? 5-20-25, 10-40-50, 4-15-24, 25-100-125 2313 0

  • 1
    4-15-24
    सही
    गलत
  • 2
    5-20-25
    सही
    गलत
  • 3
    10-40-50
    सही
    गलत
  • 4
    25-100-125
    सही
    गलत
  • उत्तर देखेंउत्तर छिपाएं
  • Workspace

उत्तर : 1. "4-15-24"
व्याख्या :

Answer: A) 4-15-24 Explanation: In each term second two numbers are multiples of the first number. But in 4-15-24 15 is not multiple of 4.

प्र: A problem is given to three students whose chances of solving it are 1/2, 1/3 and 1/4 respectively. What is the probability that the problem will be solved? 2312 0

  • 1
    1/4
    सही
    गलत
  • 2
    1/2
    सही
    गलत
  • 3
    3/4
    सही
    गलत
  • 4
    7/12
    सही
    गलत
  • उत्तर देखेंउत्तर छिपाएं
  • Workspace

उत्तर : 3. "3/4"
व्याख्या :

Answer: C) 3/4 Explanation: Let A, B, C be the respective events of solving the problem and A , B, C be the respective events of not solving the problem. Then A, B, C are independent event ∴A, B, C are independent events Now,  P(A) = 1/2 , P(B) = 1/3 and P(C)=1/4  PA=12, PB=23, PC= 34 ∴ P( none  solves the problem) = P(not A) and (not B) and (not C)                     = PA∩B∩C                    = PAPBPC         ∵ A, B, C are Independent                                          =  12×23×34                     = 14   Hence, P(the problem will be solved) = 1 - P(none solves the problem)                  = 1-14= 3/4

प्र: It takes one minute to fill 2/7 th of a vessel. What is the time taken in minutes to fill the whole of the vessel  ? 2310 0

  • 1
    1/7 min
    सही
    गलत
  • 2
    7/2 min
    सही
    गलत
  • 3
    5/7 min
    सही
    गलत
  • 4
    7/5 min
    सही
    गलत
  • उत्तर देखेंउत्तर छिपाएं
  • Workspace

उत्तर : 2. "7/2 min"
व्याख्या :

Answer: B) 7/2 min Explanation: Time taken to fill 2/7 = 1Then to fill full 1 = ?? = 1/(2/7) = 7/2 minutes.

प्र: 1, 2, 8, 33, ? Find the next number in the given series? 2310 0

  • 1
    148
    सही
    गलत
  • 2
    145
    सही
    गलत
  • 3
    143
    सही
    गलत
  • 4
    141
    सही
    गलत
  • उत्तर देखेंउत्तर छिपाएं
  • Workspace

उत्तर : 1. "148"
व्याख्या :

Answer: A) 148 Explanation: 1 x 1 + 1 x 1 = (1 + 1) = 2 2 x 2 + 2 x 2 = (4 + 4) = 8 8 x 3 + 3 x 3 = (24 + 9) = 33 33 x 4 + 4 x 4 = (132 + 16) = 148 148 x 5 + 5 x 5 = (740 + 25) = 765.

      त्रुटि की रिपोर्ट करें

    कृपया संदेश दर्ज करें
    त्रुटि रिपोर्ट सफलतापूर्वक जमा हुई

      त्रुटि की रिपोर्ट करें

    कृपया संदेश दर्ज करें
    त्रुटि रिपोर्ट सफलतापूर्वक जमा हुई

      त्रुटि की रिपोर्ट करें

    कृपया संदेश दर्ज करें
    त्रुटि रिपोर्ट सफलतापूर्वक जमा हुई

      त्रुटि की रिपोर्ट करें

    कृपया संदेश दर्ज करें
    त्रुटि रिपोर्ट सफलतापूर्वक जमा हुई

      त्रुटि की रिपोर्ट करें

    कृपया संदेश दर्ज करें
    त्रुटि रिपोर्ट सफलतापूर्वक जमा हुई

      त्रुटि की रिपोर्ट करें

    कृपया संदेश दर्ज करें
    त्रुटि रिपोर्ट सफलतापूर्वक जमा हुई

      त्रुटि की रिपोर्ट करें

    कृपया संदेश दर्ज करें
    त्रुटि रिपोर्ट सफलतापूर्वक जमा हुई

      त्रुटि की रिपोर्ट करें

    कृपया संदेश दर्ज करें
    त्रुटि रिपोर्ट सफलतापूर्वक जमा हुई